OpenStudy (anonymous):

estimate for the negative square root of 93 ?

TheSmartOne (thesmartone):

so last time we did up to 8^2 let's continue from there since 93 is a larger number 8^2 = ? 9^2 = ? 10^2 =? 11^2 = ?

OpenStudy (anonymous):

64 81 100 121

TheSmartOne (thesmartone):

and 93 fits in between which two numbers?

OpenStudy (anonymous):

81 and 100

TheSmartOne (thesmartone):

correct and which number is 93 closer to? 81 or 100?

OpenStudy (anonymous):

100

TheSmartOne (thesmartone):

correct, so what do you thing sqrt of 93 would be approximately? like would it be greater than 9.5 or less than 9.5?

OpenStudy (anonymous):

Greater?

TheSmartOne (thesmartone):

correct sqrt 93 is equal to 9.64

TheSmartOne (thesmartone):

is approximately qeual to 9.64**

OpenStudy (anonymous):

Thank you!

TheSmartOne (thesmartone):

Anytime! :)